A basketball team recently scored a total of
77 points on a combination of​ 2-point field​ goals, 3-point field​ goals, and​ 1-point foul shots.​ Altogether, the team made
44 baskets and 10 more​ 2-pointers than foul shots. How many shots of each kind were​ made?

Answer by algebrahouse.com(1659)

x = number of foul shots made
x + 10 = number of 2-pointers {they made 10 more 2-pointers than foul shots}
44 - (x + x + 10) = number of 3-pointers {they made a total of 44 baskets}
= -2x + 34

x + 2(x + 10) + 3(-2x + 34) = 77 {value of basket times number of baskets equals total points}
x + 2x + 20 - 6x + 102 = 77 {used distributive property}
-3x + 122 = 77 {combined like terms}
-3x = -45 {subtracted 122 from each side}
x = 15 {divided each side by -3}
x + 10 = 25 {substituted 15, in for x, into (x + 10)}
-2x + 34 = 4 {substituted 15, in for x, into (-2x + 34)}

15 one-point foul shots
25 two-point field goals
4 three point field goals
= poor defense

For more help like this, visit: www.algebrahouse.com